Calculate the pH of a solution containing 0.01M HF and 0.002M NaF if the Ka=7.2x10-4. Calculate the pH if 0.002M of KOH is added to the buffer.



1
Expert's answer
2021-02-17T07:45:27-0500

Molarity of "\text{HF}=0.01M"

Molarity of "\text{NAF}=0.002M"


"Ka=7.2\times 10^{-4}"


"P_{k_a}=-log(7.2\times 10^{-4})=4-0.85=3.15"


Then The "P_H" of the solution formed is given by the following formula:-

"P_H=P_{k_a}+log\dfrac{[\text{salt}]}{[\text{Acid]}}"


"=3.15+log\dfrac{[0.002]}{[0.01]}"


"=3.15+log(0.2)"


=3.15+0.301

=3.451


Now When KOH is added to the buffer KOH dissassociates to form


"KOH\rightarrow K^++OH^-"


"P_{OH}=-log([OH]^{-1})=-log(0.002)=2.7"


Also, "P_H+P_{OH}=14,P_H=14-2.7=11.3"


Hence the "P_{H}" is 11.3